OpenStudy (anonymous):

how would I simplify 10a^2 b+15bc^2 / -5b

OpenStudy (zzr0ck3r):

this is equal to (10a^2*b)/(-5b) + (15b*c^2)/(-5b) = (10/-5)*a^2*b^(1-1) + (15/-5)*b^(1-1)*c^2 = -2*a^2 + (-3)*c^2 = -2a^2-3c^2

OpenStudy (zzr0ck3r):

how i like to think of these: we devide 10a^2*b by -5b and got -2a^2 and then we devided 15bc^2 by -5b and got -3c^2 then added them together -2a^2-3c^2

OpenStudy (anonymous):

thank you!

OpenStudy (zzr0ck3r):

np